A golden rectangle is a rectangle whose length is approximately 1.6 times its width. The early Greeks thought that a rectangle with these dimensions was the most pleasing to the eye and examples of the golden rectangle are found in many early works of art. For​ example, the Parthenon in Athens contains many examples of golden rectangles. Mike Hallahan would like to plant a rectangular garden in the shape of a golden rectangle. If he has 234 feet of fencing​ available, find the dimensions of the garden.

Answer by rothauserc(4718)

let w be the width of the rectangle, then 1.6w is the length
:
2w + 2(1.6w) = 234
:
2w + 3.2w = 234
:
5.2w = 234
:
width = 45 feet
length = 45 * 1.6 = 72 feet
: